### v2

# Josée Desharnais    Ana Sokolova

Abstract
The most studied and accepted pseudometric for probabilistic processes is one based on the Kantorovich distance between distributions. It comes with many theoretical and motivating results, in particular it is the fixpoint of a given functional and defines a functor on (complete) pseudometric spaces. It is also the foundation for a categorical lifting of pseudometrics.
Other notions of behavioural pseudometrics have also been proposed, one of them ($\varepsilon$-distance) based on $\varepsilon$-bisimulation. $\varepsilon$-Distance has the advantages that it is intuitively easy to understand, it relates systems that are conceptually close (for example, an imperfect implementation is close to its specification), and it comes equipped with a natural notion of $\varepsilon$-coupling. Finally, this distance is easy to compute.
We show that $\varepsilon$-distance is also the greatest fixpoint of a functional and provides a functor. The latter is obtained by replacing the Kantorovich distance in the lifting functor with the Lévy-Prokhorov distance. In addition, we show that $\varepsilon$-couplings and $\varepsilon$-bisimulations have an appealing coalgebraic characterization.
keywords:
Lévy-Prokhorov metric, behavioural distance, epsilon-bisimulation, reactive probabilistic transition systems, discrete labelled Markov processes, coalgebraic epsilon-(bi)simulation
1 Introduction
Probabilistic systems [28, 37, 46, 6, 7, 43] and their behaviour have been the object of study of over thirty years in the area of formal verification and analysis of systems. They are used to represent uncertainty, incomplete information, as well as randomized behaviour. One important direction, in order to prove that systems behave the same, is the study of behavioural equivalences: these identify states with (exactly) the same behaviour. Behavioural equivalences are an elegant way to analyse and compare behaviour of systems, they have nice foundations in concurrency theory [40, 58] as well as elegant abstract generalizations in the theory of coalgebras [20, 47, 31, 35, 32, 34, 12, 13]. However, as was already observed in  [29], in a non-exact world, e.g., when the probabilities in the model are approximate and not exactly known, or estimated by sampling, or when small differences should not be considered the same as large differences, behavioural equivalences may be too strong.
One solution is to employ a distance, actually a pseudometric, providing a quantitative notion of how much states in a probabilistic model differ from one another, or how close they are to each other–distance zero corresponding to equivalence. The study of pseudometrics [22, 55, 56, 24, 17, 57, 53, 26, 55, 56, 51, 52] and quantitative theories [38, 39, 4, 5, 8, 41, 15, 41, 42] has been a fruitful one in the past decade(s), the references above showing only some of the original sources.
Historically, the impulse to define a behavioral distance between probabilistic processes became imperative for those with a continuous state space and continuous probability distributions, called Labelled Markov Processes (LMPs). It started with the idea of defining a real-valued logic, as proposed by Kozen [36], to express properties satisfied by processes: Taking the supremum over the differences of the values of formulas was a natural way of defining a pseudometric, see Desharnais et al. [22, 24]. The authors later on observed that this distance was the greatest fixpoint of some functional [25] and, in between and later, van Breugel and Worrell [55, 56, 57] and van Breugel et al. [53], proved that this gives a functor (monad) that is a lifting of the distribution functor (monad) using the Kantorovich metric on distributions. To our knowledge, no other such metric has been lifted to transition systems in a categorical way. Moreover, the distance is the one obtained using final coalgebra semantics. Since then, this pseudometric has been studied further in different contexts and algorithms have been proposed to compute it [49, 50, 4, 5]. The distance comes with many theoretical and motivating results, and has also been extended by Baldan et al. [8] to a generic Kantorovich-style (and dually Wasserstein) lifting of arbitrary functors parametric in certain evaluation maps—the two liftings coincide on distributions. With all these seals of approval, the Kantorovich behavioural distance has imposed itself as the one to use and study. One aspect that it lacks is an intuitive accompanying notion of approximate bisimulation.
Approximate bisimulations provide another strategy to circumvent the non-robustness of bisimulation equivalences. They have been studied in non-probabilistic systems [59, 30], and in probabilistic systems [26, 51, 27, 1, 2, 9, 48]. We focus on the notion of $\varepsilon$-bisimulation defined in [26]. One aim of such approximate bisimulations is to unite the best of both worlds: give a relational structure for reasoning about systems and at the same time define a distance. States related by an $\varepsilon$-bisimulation are at distance at most $\varepsilon$ from one another. The advantages of such a distance are its intuitive nature and its ability to relate systems that are almost the same in structure, as the next example will illustrate. Finally, the distance can be computed relatively easily.
For comparison with the Kantorovich behavioural distance, consider the imperfect channel (when $\varepsilon>0$) depicted in Figure 1. The $\varepsilon$-distance of this channel to a perfect channel
[…]
Figure 1: A simple channel that fails to send a token with probability $\varepsilon$
(one with $\varepsilon=0$) is $\varepsilon$, and it provides a simple example that imperfect implementation can be considered close to its specification. On the other hand, the Kantorovich distance (without discount) gives distance 1 to the pair of perfect and imperfect channels, which reflects the fact that the two channels have a very different behavior on the long run–see Example 2.1 for more details. Interestingly, $\varepsilon$-distance has an almost build-in notion of continuity: Two channels with close values of $\varepsilon$ are close in the $\varepsilon$-distance as well.
However, the approximate equivalences had no categorical formulation attached and it was not clear whether the associated distance shares some of the generic good properties identified for the Kantorovich metric. Our main motivation for this work was to investigate whether we can remedy this situation. Indeed, we provide the missing abstract characterization of $\varepsilon$-bisimulation in two directions: by obtaining a fixpoint characterization of the induced distance and by giving a coalgebraic view following Aczel-Mendler coalgebraic bisimilarity. The fixpoint characterization is the main result of the paper. For this, we show how to replace the Kantorovich pseudometric on distributions with the Lévy-Prokhorov pseudometric on distributions, and show that the distance obtained is indeed the $\varepsilon$-distance. It is remarkable how well the Lévy-Prokhorov distance fits into the definition of $\varepsilon$-bisimulation: We show that $\varepsilon$-distance is the greatest fixpoint of a suitable functional, and that the lifting also lifts the discrete distributions functor to a functor on pseudo-metric spaces. We also show a result that has no matching for the Kantorovich functional: any fixpoint distance of the Lévy-Prokhorov functional defines an $\varepsilon$-bisimulation. Applications of our results may take different routes: On the practical side, the distance could be computed by an iterative algorithm given a suitable fixpoint theorem. On the theoretical side, the fixpoint characterization may open the way for studying new problems, one example being $\varepsilon$-bisimulations up-to as an acceleration to computing $\varepsilon$-bisimilarity, along the lines of [11].
Our observation may also open a new way of studying other pseudometrics on probabilistic systems that would be constructed from other basic distances on distributions. More generally, the construction could be applied to a pseudometric on $FX$ for a functor $F$ on $\mathbf{Sets}$. It is yet to be seen whether the Lévy-Prokhorov lifting fits the Kantorovich / codensity - style liftings of Baldan et al. [8] which aim at generalizing the Kantorovich lifting to generic functors. If not, perhaps the lifting can provide another way of lifting functors to metric spaces.
Unlike in the case of the Kantorovich lifting, the Lévy Prokhorov lifting does not yield a monad on the category of pseudo-metric spaces, see Section 3.4. As a consequence, we could not follow the path of [53] and reproduce their results on the final coalgebra. Instead, we take a different way and describe $\varepsilon$-(bi)simulations as coalgebraic (bi)simulations, following and extending the framework of Hasuo [31, 32] and the original results of Hughes and Jacobs [33]. We only present the necessary abstraction needed to discuss $\varepsilon$-(bi)simulations on discrete time Markov chains and labelled Markov processes, generalizations of these notions are possible and we plan to elaborate on them in follow-up work.
2 The objects of interest
Our development concerns discrete labelled Markov processes (LMPs) also called reactive probabilistic transition systems. They consist of a set of states $S$ and transitions $\tau_{a}:S\to\mathcal{D}S$ labelled with actions $a$ of some set $\mathcal{A}$, where $\mathcal{D}S$ is the set of discrete sub-distributions on $S$:
$\mathcal{D}S=\{\varphi\colon S\to[0,1]\mid\sum_{s\in S}\varphi(s)\leq 1\}.$
Note that the size of the state set need not be restricted (for the sums to be defined), as the sum is defined as $\sum_{x\in S}\varphi(x)=\sup_{{\textrm{finite }}X\subseteq S}\left(\sum_{x\in X}\varphi(x)\right)$. It is easy to prove that discrete probability distributions always have a countable support, i.e., they assign non-zero probability to at most countably many elements of $S$.
For $s,t\in S$ and $a\in\mathcal{A}$, the value $\tau_{a}(s)(t)$ (also written $\tau_{a}(s,t)$) encodes the probability of jumping from $s$ to $t$ when action $a$ is taken. Figure 1 shows such a system with actions send and ack. The first transition is a subprobability, where the missing probability represents the probability of failure. If $X\subseteq S$, we write $\tau_{a}(s)(X)$ for $\sum_{t\in X}\tau_{a}(s)(t)$. Coalgebraically, LMPs, for a set of actions or labels $\mathcal{A}$, are arrows $\tau\colon S\to(\mathcal{D}S)^{\mathcal{A}}$. We write $\tau_{a}(s)$ for the subdistribution $\tau(s)(a)$. In case of just one label, LMPs are (discrete time) Markov chains, i.e., coalgebras of the functor $\mathcal{D}$. The reason for working with subdistributions is that they come equipped with a non-trivial pointwise order. Bisimilarity of LMPs is the classical Larsen-Skou [37] bisimilarity.
We denote by $\mathbf{PMet_{1}}$ the category of $1$-bounded pseudometric spaces. It has as objects pseudometric spaces, which are pairs $\langle X,d\rangle$ of a set $X$ and a function $d\colon X\times X\to[0,1]$ which satisfies $d(x,x)=0$, is symmetric: $d(x,y)=d(y,x)$ and satisfies the triangular inequality: $d(x,y)+d(y,z)\geq d(x,z)$ for all $x,y,z\in X$. As arrows, $\mathbf{PMet_{1}}$ has non-expansive maps $f\colon\langle X,d_{X}\rangle\to\langle Y,d_{Y}\rangle$, i.e., functions $f\colon X\to Y$ satisfying $d_{Y}(f(x_{1}),f(x_{2}))\leq d_{X}(x_{1},x_{2})$. A metric or distance is a pseudometric that additionally satisfies $d(x,y)=0$ iff $x=y$. Of particular interest for LMPs are pseudometrics whose kernel is bisimilarity, that is, $d(x,y)=0$ iff $x$ and $y$ are bisimilar. As we only deal with pseudometrics here, we often use the word distance or metric as a shorthand for pseudometric. Following [57, 53], we will use the opposite pointwise order [^1] The only reason for this choice is to obtain the distance as the greatest, rather than least, fixpoint – just like bisimilarity is the greatest fixpoint of a suitable functional. on distances defined by:
$d_{1}\sqsubseteq d_{2}\quad\Longleftrightarrow\quad\forall x,y\in X:d_{1}(x,y)\geq d_{2}(x,y).$
However, whenever it is clearer, we may use the direct pointwise order too: $d_{1}\geq d_{2}\Leftrightarrow d_{1}\sqsubseteq d_{2}$.
2.1 The Kantorovich pseudometric
The observation that bisimulation is too strong in the context of probabilistic systems has led to the idea of defining a pseudometric that would give zero distance to bisimilar states. Initially, this pseudometric was defined using a real-valued logic. A set $\cal F$ of functionals were defined from states of LMPs to $[0,1]$, with the following syntax, mimicking logical formulas [24]:
$f:=1\mid\inf(f_{1},f_{2})\mid 1-f\mid f\varominus q\mid\langle a\rangle\,f\mbox{ with }q\in[0,1]\cap\mathbb{Q}.$
We omit the semantics, but the next example will give a taste of it. A distance emerged naturally as
$d_{K}(s,t)=\sup_{f\in\cal F}\big{|}f(s)-f(t)\big{|}.$
Example 2.1.
As an example, we can look back at the channel of Figure 1. The maximum difference over functionals between $c_{\varepsilon}$ and $c_{\gamma}$, with $\varepsilon,\gamma>0$ is over the functional $\langle\textit{send\,}\rangle\langle\textit{ack}\,\rangle 1$, which evaluates to $1-\varepsilon$ and $1-\gamma$, respectively, and thus yields the distance $d_{K}(c_{\varepsilon},c_{\gamma})=|\varepsilon-\gamma|$. Between $c_{\varepsilon}$ and $c_{0}$, the functional $(\langle send\rangle\langle ack\rangle)^{n}1$ evaluates to $(1-\varepsilon)^{n}$ on $c_{\varepsilon}$, and to 1 on $c_{0}$. So the supremum results in $d_{K}(c_{\varepsilon},c_{0})=1$.
Later on, it was proven that the functionals could be any non-expansive maps and that this distance is the distance obtained via a final coalgebra construction for a functor (monad) that is a lifting of the Distribution functor (monad) to pseudometric spaces using the Kantorovich metric, by van Breugel and Worrell [57] and van Breugel et al. [53, 54].
This distance is a fixpoint of the functional on distances on the states of an LMP $\tau:S\to(\mathcal{D}S)^{\mathcal{A}}$
$\Delta_{K}(d)(s_{0},s_{1})=\sup_{a\in\mathcal{A}}\,\delta_{K}^{d}(\tau_{a}(s_{0}),\tau_{a}(s_{1}))$ (1)
where $\delta_{K}^{d}$ is the standard well-known Kantorovich distance between two distributions. We do not need its definition here, the interested reader can find it in, e.g., [53].
The distance on distributions is just a parameter in Equation (1); changing the Kantorovich distance $\delta_{K}^{d}$ to another distance on distributions and looking for a fixpoint will give us a new distance on states. In fact we are interested in the question formulated the other way around: find a distance on distributions for which the fixpoint is the $\varepsilon$-distance $d^{*}$, defined in the next section. The functional $\Delta_{K}$ was discovered after the behavioural distance $d_{K}$ was introduced as well. However, it could have been done the other way around, choosing a distance on distributions, and looking for the resulting distance on states, similarly as how it is done by Baldan et al. [8]. We discuss this in more detail again in Section 3 below.
2.2 $\varepsilon$-Bisimulation
Bisimulation being too strong for the comparison of states in quantitative systems, has also led to relaxing the definition of simulation and bisimulation itself to approximate relations. As mentioned in the introduction, there are a few of those definitions of approximate bisimulations, but we are interested in the following.
Definition 2.2 ([26]).
Let $\tau:S\to(\mathcal{D}S)^{\mathcal{A}}$ be an LMP and let $\varepsilon\in[0,1]$. A relation $R\subseteq S\times S$ is an $\varepsilon$-simulation if whenever $sRt$, then for all $a\in\mathcal{A}$, for all $X\subseteq S$
$\tau_{a}(s)(X)\leq\tau_{a}(t)({R}(X))+\varepsilon,\qquad\mbox{ where }R(X)=\{y\mid\exists x\in X:xRy\}.$
If $R$ is symmetric, it is an $\varepsilon$-bisimulation.
A state $s$ is $\varepsilon$-simulated by state $t$, written $s\prec_{\varepsilon}t$, if $sRt$ for some $\varepsilon$-simulation $R$.
If $sRt$ for $R$ $\varepsilon$-bisimulation, we write $s\sim_{\varepsilon}t$ and we say that $s$ and $t$ are $\varepsilon$-bisimilar.
As expected, ordinary (bi)simulation on LMPs [37, 10] is simply 00-(bi)simulation. This definition has an extension to nondeterministic and probabilistic finite systems [51].
The operation $R(X)$ on a set $X$ is what restricts this work to discrete distributions. In general, if $X$ is measurable, we may not have $R(X)$ measurable. Working in analytic spaces, as was done before for LMPs [23] may solve this issue, but we leave it for future work. In particular, one could ask $R(X)$ to be measurable whenever $X$ is, as done in [9], but many results are not proven in that case, like the logical characterisation, or Theorem 2.5 below.
Example 2.3.
Consider the following example, with $\gamma\in(0,1]$.
The relation $R=\{(s,t),(s_{1},t_{1})\}$ is an $\varepsilon$-simulation for $\varepsilon=\gamma$, and so is $R\cup R^{-1}$. Hence, $s\prec_{{\gamma}}t$, $t\prec_{{\gamma}}s$ and $s\sim_{{\gamma}}t$. However, the situation is different for $\varepsilon<\gamma$: for example, taking $\varepsilon=0$, we observe that $s\prec_{0}t$ (with the relation $R\cup\{(s_{2},t_{1})\}$). However, $t\not\prec_{0}s$, because a relation $R^{\prime}$ relating $t$ and $s$ would have to also relate $t_{1}$ to $s_{1}$ and to $s_{2}$, if $\gamma<1$, and $t_{1}$ to $s_{2}$ if $\gamma=1$. Indeed, taking $X=\{t_{1}\}$, we need $\tau_{a}(t)(X)\leq\tau_{a}(s)(R^{\prime}(X))+0$, but for this to be the case, we would need $\tau_{a}(s)(R^{\prime}(X))=1$ and hence $R^{\prime}$ must include the pairs $(t_{1},s_{1})$ and $(t_{1},s_{2})$ if $\gamma<1$, and otherwise must include $(t_{1},s_{2})$ if $\gamma=1$. So, in any case $(t_{1},s_{2})\in R^{\prime}$, but the pair $(t_{1},s_{2})$ cannot be related by any $\varepsilon$-simulation relation $R^{\prime}$ for $\varepsilon<1$ (which is the case here as $\varepsilon=0$), as, for example, taking again $X=\{t_{1}\}$, we would have $\tau_{a}(t_{1})(X)=1$ but $\tau_{a}(s_{2})(R^{\prime}(X))=0$. Hence, also $s\not\sim_{0}t$.
In contrast to the case of 0-bisimulation, two-way $\varepsilon$-simulation is not $\varepsilon$-bisimulation [26, 51]. From the notion of $\varepsilon$-bisimilarity, a pseudometric arises naturally by taking the infimum.
Definition 2.4 (The $\varepsilon$-distance $d^{*}$ [26]).
Let $\tau:S\to(\mathcal{D}S)^{\mathcal{A}}$ be an LMP. The pseudometric $d^{*}$ on $S\times S$, called $\varepsilon$-distance, is defined as follows:
$\begin{array}[]{lcl}d^{*}:&S\times S&\rightarrow[0,1]\\ &(s,t)&\mapsto\inf\ \left\{\varepsilon\in[0,1]\mid s\sim_{\varepsilon}t\right\}.\end{array}$
The function $d^{*}$ is a pseudometric on the states of the considered LMP: the triangle inequality comes from the well-known fact, see e.g. [26], that $s\sim_{\varepsilon_{1}}u$ and $u\sim_{\varepsilon_{2}}t$ imply $s\sim_{\varepsilon_{1}+\varepsilon_{2}}t$. (The same property holds for $\prec_{\varepsilon}$ as well.) Indeed, $\prec_{\varepsilon}$ and $\sim_{\varepsilon}$ are not transitive. Instead, these relations are entourages that form a uniform structure, or uniformity, cf. [14].
The kernel of $d^{*}$, that is, the set of pairs at zero distance, is $\sim_{0}$, the bisimilarity relation on LMPs.
Theorem 2.5 ([26]).
$d^{*}(s,t)=0\Leftrightarrow s\sim_{0}t$.
One of the nice properties of this distance is its relatively easy computability, both regarding complexity and regarding computation "by hand". It suffices to define an $\varepsilon$-bisimulation to obtain an upper bound to the distance between states. In Example 2.3, we have $d^{*}(s,t)=\gamma$, and, indeed, we easily see that the probabilities of $s$ and $t$ seen as processes are within $\gamma$. Also here, the Kantorovich distance between these states is 1. One way to see that is by using the functional $\langle a\rangle^{n}1$, which evaluates to $(1-\varepsilon)^{n}$ on $s$, and to 1 on $t$. So $d_{K}(s,t)=1$. Similarly, as already noted in the introduction, for the channel in Figure 1, we can prove that $d^{*}(c_{\varepsilon},c_{\gamma})=|\varepsilon-\gamma|=d_{K}(c_{\varepsilon},c_{\gamma})$, for $\varepsilon,\gamma>0$, but $d^{*}(c_{\varepsilon},c_{0})=\varepsilon\neq d_{K}(c_{\varepsilon},c_{0})$.
3 The Lévy-Prokhorov distance lifting
We will now explain in details what we think should be a behavioral metric on LMPs. These details may not be new, or may appear obvious to the expert reader, but we find it useful to spell them out here. When looking for a distance $d^{\dagger}$ on the states of an LMP, one observes that states are both targets of distributions and (a set of) distributions themselves (since they are defined by their outgoing transition distributions): that is, in Example 2.3, the states $s$ and $t$ can be viewed as distributions over the set $\{s_{1},s_{2},t_{1}\}$. As there are already a few distances on distributions that were studied outside computer science and concurrency theory, they give a starting point for distances on states: we could say, given such a distance $\delta$:
$d^{\dagger}(s,t):=\sup_{a\in\mathcal{A}}\,\delta(\tau_{a}(s),\tau_{a}(t)),\qquad\mbox{ for $s,t\in S$.}$ (2)
We will use the letter $d$ for distances on states, and the symbol $\delta$ for distances on distributions. Analysing the options of this equation in full generality is outside the scope of the current paper [^2] If the reader knows such works, we would be happy to list them here., but some distances, such as $\delta_{K}^{d}$ (and the one we introduce below $\delta^{d}_{LP}$) have a particularity. They are defined using a parameter $d$, a basic distance on the space where the distributions are defined. This is desirable for concurrency theory because distributions on states that are different but close (or even bisimilar!) should also be close: so a starting metric $d$ on the state space is of key importance to account for these similarities between states/processes. Note that a few papers [30, 18] start from metrics on the states that account for extra information (e.g. a distance on the labels/observations that are attached to states)–but this goes beyond the behaviour of states that we want to capture here (although they constitute an interesting line of work to extend our method).
A needed property for $d^{\dagger}$ to be a behavioural distance, as also pointed out in [8] for $\delta_{K}^{d}$, is that it accords with the starting distance $d$ used in $\delta^{d}$, as follows.
$d^{\dagger}(s,t)=\sup_{a\in\mathcal{A}}\,\delta^{d^{\dagger}}(\tau_{a}(s),\tau_{a}(t)),\qquad\mbox{ for $s,t\in S$.}$ (3)
This says that the distance between states viewed as simple members of the space (the left hand part of the equality), is the same as their distance when viewed as processes, that is, their outgoing transition distributions. So this really says that $d^{\dagger}$ treats states according to their behaviour. Technically, this is saying that we are looking for a distance fixpoint $d^{\dagger}$ of the functional
$\Delta(d)(s,t):=\sup_{a\in\mathcal{A}}\,\delta^{d}(\tau_{a}(s),\tau_{a}(t)),\qquad\mbox{ for $s,t\in S$.}$ (4)
Of course, this is exactly Equation (1), for $\delta^{d}:=\delta_{K}^{d}$, the fixpoint property of the Kantorovich behavioural distance, observed by [25, 57], and explicitly proven in [54]. One of our goals in this work was to find a similar functional that would have the $\varepsilon$-distance $d^{*}$ as its greatest fixed point. That is, we are seeking a suitable distance $\delta^{d}$ on distributions.
Remark 3.1.
Another property that one might consider natural and one might expect from such a distance $\delta^{d}$ is that the distance between the Dirac distributions on states is the same as the starting distance between them. That is, for a distance $d$ on states, one might expect
$d(s,t)=\delta^{d}(1_{s},1_{t}),\qquad\mbox{ for $s,t\in S$,}$ (5)
where we write $1_{s}$ for the Dirac distribution on $s\in S$. This is an interesting property of the distance $\delta^{d}$ — it is a stronger version of the non-expansiveness of the unit of the distribution monad, showing that the unit is an isometry, and we will return to it in Section 3.4 below where we show that it holds for the Lévy-Prokhorov distance, for any starting distance $d$ on states. It also holds for the (undiscounted) Kantorovich distance—this is easy to prove using the Wasserstein formulation of the Kantorovich distance, whereas it does not hold for the total variation distance [42]. Therefore, this property is neither unique to the Lévy-Prokhorov distance that we are interested in, nor it is “behavioural” in the sense that we explained so far: it does not take into account the transition structure $\tau$ of the states. For a somewhat behavioural explanation, note that combining it with Equation (2), this condition implies that a pair of states $s^{\prime},t^{\prime}$ each having a single outgoing transition as a Dirac on state $s$ and $t$ respectively, would get the same distance as $s$ and $t$. In particular, such a distance does not discount the future.
3.1 The Lévy-Prokhorov distance on distributions
While examining equation (3) and trying to make $d^{*}$ fit into it as $d^{\dagger}$ we came up with the following distance on distributions. Only afterwards, we discovered that the distance was actually known as the Lévy-Prokhorov lifting of a distance $d$ to distributions. In this section we introduce the distance and provide an example that illustrates it.
Definition 3.2 (Lévy-Prokhorov distance [44]).
Let $\langle S,d\rangle$ be a metric space ; we endow $\mathcal{D}S$ with the pseudo-metric $\delta^{d}_{LP}:\mathcal{D}S\times\mathcal{D}S\to[0,1]$, defined as
$\delta^{d}_{LP}(\mu_{0},\mu_{1})=\inf\{\varepsilon\mid\forall X\subseteq S:\mu_{i}(X)\leq\mu_{1-i}(X^{d}_{\varepsilon})+\varepsilon,\mbox{ for }i=0,1\},$
where $X^{d}_{\varepsilon}=\{y\mid\exists x\in X:d(x,y)<\varepsilon\}$.
We call $\delta^{d}_{LP}$ the Lévy-Prokhorov (LP, for short) distance.
In the following example we define simple probability measures that help illustrate the need for the extra “$+\,\varepsilon$” in this definition and the need for the “$\varepsilon$ ball around $X$”. At first sight, $\delta^{d}_{LP}$ looks very much like the total variation distance, but this “$\varepsilon$ ball around $X$” makes it very different.
Example 3.3.
Consider the set of states $S=\{x_{\gamma}\mid\gamma\in[0,1]\}$ and a distance on these states given by $d(x_{\gamma},x_{\xi})=|\gamma-\xi|$, for $\gamma,\xi\in[0,1]$. In a probabilistic transition system, one could imagine that $x_{\gamma}$ has an $a$-loop to itself with probability $1-\gamma$, $\gamma\in[0,1]$, as depicted below (adding transitions out of the states to help see their differences as processes in a transition system). We now define a family of distributions on these states, and we picture them below as states with an outgoing transition without label (one could imagine an $a$-label). Let $\nu_{\gamma}=\gamma 1_{x_{1}}+(1-\gamma)1_{x_{\gamma}}$, for $\gamma\in[0,\frac{1}{2})$, with $1_{x}$ the Dirac measure on $x$. The distributions $\nu_{\gamma}$ and $\nu_{0}$ are illustrated below. For a fixed $\gamma$, these two distributions are actually non-zero on a three-state space $S^{\prime}=\{x_{0},x_{1},x_{\gamma}\}$
We show that $\delta^{d}_{LP}(\nu_{\gamma},\nu_{0})=\gamma$. Let $1-\gamma>\varepsilon>\gamma$. We show that all such $\varepsilon$ satisfy the inequalities in the definition of $\delta^{d}_{LP}$ and any $\varepsilon\leq\gamma$ does not, so the distance, being the infimum over all such values, is indeed $\gamma$. In particular, we need to check that for all $X\subseteq S$, $\nu_{\gamma}(X)\leq\nu_{0}(X_{\varepsilon}^{d})+\varepsilon$ and $\nu_{0}(X)\leq\nu_{\gamma}(X_{\varepsilon}^{d})+\varepsilon$. The cases $X=S$ and $X=\emptyset$ are trivial, since the distributions are full. For the sets $X$ of size two, we have
$\begin{array}[]{lllll}\mbox{For }X:=\{x_{0},x_{1}\}:&\nu_{\gamma}(X)=\gamma&\!\!\leq 1+\varepsilon&=\nu_{0}(S)+\varepsilon&=\nu_{0}(X^{d}_{\varepsilon})+\varepsilon\\ &\nu_{0}(X)=1&\!\!\leq 1+\varepsilon&=\nu_{\gamma}(S)+\varepsilon&=\nu_{\gamma}(X^{d}_{\varepsilon})+\varepsilon\\ \mbox{For }X:=\{x_{1},x_{\gamma}\}:&\nu_{\gamma}(X)=1&\!\!\leq 1+\varepsilon&=\nu_{0}(S)+\varepsilon&=\nu_{0}(X^{d}_{\varepsilon})+\varepsilon\\ &\nu_{0}(X)=0&\!\!\leq 1+\varepsilon&=\nu_{\gamma}(S)+\varepsilon&=\nu_{\gamma}(X^{d}_{\varepsilon})+\varepsilon\\ \mbox{For }X:=\{x_{0},x_{\gamma}\}:&\nu_{\gamma}(X)=1-\gamma&\!\!\leq 1+\varepsilon&=\nu_{0}(\{x_{0},x_{\gamma}\})+\varepsilon&=\nu_{0}(X^{d}_{\varepsilon})+\varepsilon\\ &\nu_{0}(X)=1&\!\!\leq(1-\gamma)+\varepsilon&=\nu_{\gamma}(\{x_{0},x_{\gamma}\})+\varepsilon&=\nu_{\gamma}(X^{d}_{\varepsilon})+\varepsilon\\ \end{array}$
The last inequality is satisfied thanks to the room given by the “$+\varepsilon$”. For sets of size one, omitting the checks where the probability is zero on the left-hand side of the inequality:
$\begin{array}[]{lllll}\mbox{For }X:=\{x_{1}\}:&\nu_{\gamma}(X)=\gamma&\!\!\leq 0+\varepsilon&=\nu_{0}(\{x_{1}\})+\varepsilon&=\nu_{0}(X^{d}_{\varepsilon})+\varepsilon\\ \mbox{For }X:=\{x_{\gamma}\}:&\nu_{\gamma}(X)=1-\gamma&\!\!\leq 1+\varepsilon&=\nu_{0}(\{x_{\gamma},x_{0}\})+\varepsilon&=\nu_{0}(X^{d}_{\varepsilon})+\varepsilon\\ \mbox{For }X:=\{x_{0}\}:&\nu_{0}(X)=1&\!\!\leq 1-\gamma+\varepsilon&=\nu_{\gamma}(\{x_{\gamma},x_{0}\})+\varepsilon&=\nu_{\gamma}(X^{d}_{\varepsilon})+\varepsilon.\end{array}$
In the one but last inequality, it is crucial that $x_{\gamma}$ be within $\varepsilon$ of $x_{0}$, which includes $x_{0}$ in the $\varepsilon$ ball around $x_{\gamma}$. With $X$ in place of $X^{d}_{\varepsilon}$, the inequality would not be satisfied, and not even with $X^{d}_{0}$, a set where states at distance zero are included (like bisimilar states–of which there are none here), because we would obtain:
$\nu_{\gamma}(\{x_{\gamma}\})=1-\varepsilon\not\leq 0+\varepsilon=\nu_{0}(\{x_{\gamma}\}^{d}_{0})+\varepsilon.$
The last inequality also illustrates this, and also the need for the “$+\varepsilon$”, as the $\varepsilon$-ball itself is not enough for the inequality to be satisfied. Moreover, for a pair of distributions $\nu_{\gamma},\nu_{\xi}$ in our family one can similarly prove that $\delta^{d}_{LP}(\nu_{\gamma},\nu_{\xi})=|\gamma-\xi|$.
Finally, we note that for $\varepsilon\leq\gamma$ we have $X_{\varepsilon}^{d}=X$ for any $X\subseteq S$ and hence, e.g., for $X=\{x_{\gamma}\}$, as already noted above we have $\nu_{\gamma}(X)=1-\gamma\not\leq 0+\varepsilon=\nu_{0}(X^{d}_{\varepsilon})+\varepsilon$ as $\varepsilon\leq\gamma<\frac{1}{2}$.
One can notice that when viewed as processes, $\nu_{\gamma}$ and $\nu_{0}$ are the same as $s$ and $t$ of Example 2.3. Technically, it is rather $\tau_{a}(s)$ and $\tau_{a}(t)$ of course.
Once one has seen the Lévy-Prokhorov distance on distributions it seems not surprising that it has some link with the $\varepsilon$-distance. However, our surprise was the other way around, when we realized that the distance on distributions $\delta^{d}$ that we had to concoct to make $d^{*}$ a fixed point of $\Delta$ in Equation (4) already existed.
3.2 The Lévy-Prokhorov distance on LMPs
We are now ready to define the functional on distances over LMPs that we were looking for. We expect a fixpoint of this functional to be a behavioral metric, but more importantly we expect $d^{*}$ to be the greatest fixpoint of it.
Definition 3.4.
Let $\langle S,d\rangle$ be a pseudometric space, on which an LMP $\tau:S\to(\mathcal{D}S)^{\mathcal{A}}$ is defined. Let $s_{0}$ and $s_{1}$ be states. We define the functional $\Delta_{LP}$ as
$\Delta_{LP}(d)(s_{0},s_{1})$ $=\sup_{a\in\mathcal{A}}\delta_{LP}^{d}(\tau_{a}(s_{0}),\tau_{a}(s_{1}))$
$=\sup_{a\in\mathcal{A}}\inf\{\varepsilon\mid(\forall A\subseteq S:\tau_{a}(s_{i})(A)\leq\tau_{a}(s_{1-i})(A^{d}_{\varepsilon})+\varepsilon,\quad i=0,1)\}.$
Hence, for a fixed pseudometric space $\langle S,d\rangle$, $\Delta_{LP}$ applied to $d$ gives a new pseudometric on $S$. Clearly, this can be seen as a functional on $\mathbf{PMet_{1}}$, mapping $\langle S,d\rangle$ to $(S,\Delta_{LP}(d))$.
Before proving that $d^{*}$ is a fixpoint, we show how the supremum over actions can safely be replaced by inserting a universal quantification on actions inside the set, as will be useful in Proposition 3.16.
Lemma 3.5.
Let $\langle S,d\rangle$ be a pseudometric space, on which an LMP $\tau:S\to(\mathcal{D}S)^{\mathcal{A}}$ is defined. Let $s_{0}$ and $s_{1}$ be states. Then
$\Delta_{LP}(d)(s_{0},s_{1})$ $=\inf\{\varepsilon\mid(\forall a\in\mathcal{A},\forall A\subseteq S:\tau_{a}(s_{i})(A)\leq\tau_{a}(s_{1-i})(A^{d}_{\varepsilon})+\varepsilon,\quad i=0,1)\}.$
Proof 3.6.
For $a\in\mathcal{A}$ and $\varepsilon>0$, we define the following predicate:
$P(a,\varepsilon):=(\forall A\subseteq S:\tau_{a}(s_{i})(A)\leq\tau_{a}(s_{1-i})(A^{d}_{\varepsilon})+\varepsilon,\quad i=0,1).$
This family of predicates satisfies
$\mbox{if }\gamma>\varepsilon,P(a,\varepsilon)\Rightarrow P(a,\gamma).$ (6)
We now prove that
$\sup_{a\in\mathcal{A}}\inf\{\varepsilon\mid P(a,\varepsilon)\}=\inf\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}$
  • $\leq$:
    The following sequence of arguments shows this inequality.
    $\{\varepsilon\mid P(a_{0},\varepsilon)\}$$\supseteq\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}$ for all $\mbox{for all }a_{0}\in\mathcal{A},\mbox{ so}$
    $\inf\{\varepsilon\mid P(a_{0},\varepsilon)\}$ $\leq\inf\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}$ for all $\mbox{for all }a_{0}\in\mathcal{A},\mbox{ and so}$
    $\sup_{a\in\mathcal{A}}\inf\{\varepsilon\mid P(a,\varepsilon)\}$ $\leq\inf\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}.$
  • $\geq$:
    Let $\alpha=\sup_{a\in\mathcal{A}}\inf\{\varepsilon\mid P(a,\varepsilon)\}.$ Then for all $a\in\mathcal{A}$, $\alpha\geq\inf\{\varepsilon\mid P(a,\varepsilon)\}.$ Let $\gamma>\alpha$. Then we have
[…]
    so $\quad\gamma\in\cap_{a\in\mathcal{A}}\{\varepsilon\mid P(a,\varepsilon)\}$
    so $\quad\gamma\in\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}$
    so $\quad\gamma\geq\inf\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}.$
    This implies $\alpha\geq\inf\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}$ as well. Otherwise, suppose $\alpha<\iota:=\inf\{\varepsilon\mid(\forall a\in\mathcal{A}:P(a,\varepsilon))\}$. Then $\alpha<(\alpha+\iota)/2<\iota$, so the result above with $\gamma=(\alpha+\iota)/2$ yields $(\alpha+\iota)/2\geq\iota$, a contradiction.
One remarkable property of fixpoints of this functional is that they define $\varepsilon$-bisimulations.
Theorem 3.7.
Let $\tau:S\to(\mathcal{D}S)^{\mathcal{A}}$ be an LMP with $\langle S,d\rangle$ a pseudometric space. If $d$ is a fixpoint of $\Delta_{LP}$, then for any $\varepsilon>0$, $R_{\varepsilon}:=\{(s,t)\in S\times S\mid d(s,t)<\varepsilon\}$, is an $\varepsilon$-bisimulation.
Proof 3.8.
Let $d=\Delta_{LP}(d)$ be a fixpoint. Let $R_{\varepsilon}=\{(s,t)\mid d(s,t)<\varepsilon\}$. We show it is an $\varepsilon$-bisimulation. Let $s_{0}R_{\varepsilon}s_{1}$, i.e., $d(s_{0},s_{1})<\varepsilon$, and also $\Delta_{LP}(d)(s_{0},s_{1})<\varepsilon$, because $d$ is a fixpoint. Let $a\in\mathcal{A}$; then $\delta_{LP}^{d}(\tau_{a}(s_{0}),\tau_{a}(s_{1}))<\varepsilon$. So there is some $\gamma<\varepsilon$ such that for $i=1,2$ and $X\subseteq S$
$\tau_{a}(s_{i})(X)$ $\leq\tau_{a}(s_{1-i})(X^{d}_{\gamma})+\gamma$ by definition of $\delta_{LP}$
$\leq\tau_{a}(s_{1-i})(X^{d}_{\varepsilon})+\varepsilon$ because $\mbox{because }X^{d}_{\gamma}\subseteq X^{d}_{\varepsilon}\mbox{ and }\gamma\leq\varepsilon.$
$=\tau_{a}(s_{1-i})(R_{\varepsilon}(X))+\varepsilon$ by definition of $\mbox{by definition of }R_{\varepsilon}(X).$
So $R_{\varepsilon}$ is an $\varepsilon$-bisimulation, as wanted.
This theorem is nice in itself but it also has an important corollary, that $d^{*}$ is greater than or equal to all fixpoints of $\Delta_{LP}$.
Corollary 3.9.
$d^{*}$ is greater than or equal to all fixpoints, i.e.,
$\Delta_{LP}(d)=d\quad\Longrightarrow\quad d\sqsubseteq d^{*}.$
Proof 3.10.
Let $d$ be a fixpoint of $\Delta_{LP}$ and let $d(s,t)=\varepsilon$. By the previous theorem, $s\sim_{\gamma}t$ for all $\gamma>\varepsilon$. Since $d^{*}(s,t)=\inf\{e\mid s\sim_{e}t\}$, we obtain $d^{*}(s,t)\leq\varepsilon=d(s,t)$, as wanted.
Example 3.11.
One other fixpoint of $\Delta_{LP}$ is the exact bisimulation distance, which we write $d_{\sim}$. It is equal to 0 if the states are bisimilar, otherwise it is 1. We first prove that bisimilar states $s,t$ satisfy $\Delta_{LP}(d_{\sim})(s,t)=0$. Consider $R$ to be bisimilarity. Because $R$ is a bisimulation, we have that for all $s_{0}Rs_{1}$, $a\in\mathcal{A}$ and $X\subseteq S$,
$\tau_{a}(s_{0})(X)\leq\tau_{a}(s_{1})(R(X))=\tau_{a}(s_{1})(X^{d_{\sim}}_{\gamma}),$
for any $\gamma\in(0,1)$, because $X^{d_{\sim}}_{\gamma}$ is just the smallest $R$-closed set that contains $X$. A symmetric argument applies with 0 and 1 interchanged. Consequently, $\Delta_{LP}(d_{\sim})(s_{0},s_{1})=0$, as wanted.
Conversely, if $\Delta_{LP}(d_{\sim})(s_{0},s_{1})=0$, then there is a sequence of $\gamma_{n}\in(0,1)$ that converges to zero, for which we have
$\tau_{a}(s_{0})(X)\leq\tau_{a}(s_{1})(X^{d_{\sim}}_{\gamma_{n}})+\gamma_{n}=\tau_{a}(s_{1})(R(X))+\gamma_{n}.$
Since the property is monotone, recall also (6), for all $\gamma\in(0,1)$ we have
$\tau_{a}(s_{0})(X)\leq\tau_{a}(s_{1})(X^{d_{\sim}}_{\gamma})+\gamma=\tau_{a}(s_{1})(R(X))+\gamma.$
So $\tau_{a}(s_{0})(X)\leq\tau_{a}(s_{1})(R(X))$, as wanted.
We prove that $\Delta_{LP}$ is monotonic, after the following simple lemma.
Lemma 3.12.
If $d_{2}\sqsubseteq d_{1}$, we have $X^{d_{2}}_{\varepsilon}\subseteq X^{d_{1}}_{\varepsilon}$.
Proof 3.13.
Assume $d_{1}\leq d_{2}$. Let $y\in X^{d_{2}}_{\varepsilon}$. Then there is some $x\in X$ such that $d_{2}(x,y)<\varepsilon$. This gives $d_{1}(x,y)\leq d_{2}(x,y)<\varepsilon$. So $y\in X^{d_{1}}_{\varepsilon}$.
Proposition 3.14.
$\Delta_{LP}$ is monotonic.
Proof 3.15.
Let $d_{1}\leq d_{2}$. Let $s_{0}$, $s_{1}$ be states, let $a\in\mathcal{A}$ and define the two sets
$E_{1}$ =$=\{\varepsilon\mid(\forall A\subseteq S:\tau_{a}(s_{i},A)\leq\tau_{a}(s_{1-i},A^{d_{1}}_{\varepsilon})+\varepsilon,\quad i=0,1)\}$
$\mbox{and}\quad E_{2}$ =$=\{\varepsilon\mid(\forall A\subseteq S:\tau_{a}(s_{i},A)\leq\tau_{a}(s_{1-i},A^{d_{2}}_{\varepsilon})+\varepsilon,\quad i=0,1)\}.$
Then $E_{2}\subseteq E_{1}$. Indeed, let $\varepsilon\in E_{2}$ and $A\subseteq S$. Then $\tau_{a}(s_{i},A)\leq\tau_{a}(s_{1-i},A^{d_{2}}_{\varepsilon})+\varepsilon\leq\tau_{a}(s_{1-i},A^{d_{1}}_{\varepsilon})+\varepsilon$, because $A^{d_{2}}_{\varepsilon}\subseteq A^{d_{1}}_{\varepsilon}$ by Lemma 3.12. So $\inf E_{1}\leq\inf E_{2}$ and hence $\Delta_{LP}(d_{1})=\sup_{a\in\mathcal{A}}\inf E_{1}\leq\sup_{a\in\mathcal{A}}\inf E_{2}=\Delta_{LP}(d_{2})$, as wanted.
Here is the long-awaited result, that the distance defined by $\varepsilon$-bisimulation is a fixpoint of $\Delta_{LP}$.
Proposition 3.16.
$d^{*}$ is a fixpoint of $\Delta_{LP}$, and hence it is the greatest fixpoint.
Proof 3.17.
Let $s_{0}$, $s_{1}\in S$. We want that $\Delta_{LP}(d^{*})(s_{0},s_{1})=d^{*}(s_{0},s_{1})$. Consider the following two sets
$E^{\Delta}$ =$=\{\varepsilon\mid\forall a\in\mathcal{A}:(\forall X\subseteq S:\tau_{a}(s_{i},X)\leq\tau_{a}(s_{1-i},X^{d^{*}}_{\varepsilon})+\varepsilon,\quad i=0,1)\}$
$\mbox{and}\quad E^{*}$ =$=\{\varepsilon\mid s_{0}\sim_{\varepsilon}s_{1}\}$
=$=\{\varepsilon\mid\exists R\mbox{ an $\varepsilon$-bisimulation s.t. $s_{0}Rs_{1}$, that is, for all $(t_{0},t_{1})\in R$, we have}$
 $\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\quad\forall a\in\mathcal{A}:(\forall X\subseteq S:\tau_{a}(t_{i},X)\leq\tau_{a}(t_{1-i},R(X))+\varepsilon,\quad i=0,1)\}.$
Then $\Delta(d^{*})(s_{0},s_{1})=\inf E^{\Delta}$ by Lemma 3.5 and $d^{*}(s_{0},s_{1})=\inf E^{*}$. We prove that $\inf E^{\Delta}=\inf E^{*}$:
  • $\geq$:
    We prove that $E^{\Delta}\subseteq E^{*}$. Let $\varepsilon\in E^{\Delta}$. Let $R_{\varepsilon}=\{(x,y)\mid x\sim_{\varepsilon}y\}$, that is, the biggest $\varepsilon$-bisimulation. We need to prove that $s_{0}\sim_{\varepsilon}s_{1}$. For that we prove that $R_{\varepsilon}^{+}:=R_{\varepsilon}\cup\{(s_{0},s_{1})\}$ is an $\varepsilon$-bisimulation. Let $a\in\mathcal{A}$ and $X\subseteq S$. Then $X_{\varepsilon}^{d^{*}}\subseteq R^{+}_{\varepsilon}(X)$. Indeed, let $y\in X_{\varepsilon}^{d^{*}}$ then $d^{*}(y,x)<\varepsilon$ for some $x\in X$. So, because $d^{*}(y,x)=\inf\{e\mid y\sim_{e}x\}$, we have $y\sim_{\gamma}x$ for some $\gamma<\varepsilon$, and hence $y\sim_{\varepsilon}x$ as every $\gamma$-bisimulation is an $\varepsilon$-bisimulation. Hence $y\in R^{+}_{\varepsilon}(X).$ We only need to prove the bisimulation condition for the pair $(s_{0},s_{1})$. Then, for $i=0,1$ we have
    $\tau_{a}(s_{i},X)$ $\leq\tau_{a}(s_{1-i},X_{\varepsilon}^{d^{*}})+\varepsilon$ by choice of $\mbox{by choice of }\varepsilon$
    $\leq\tau_{a}(s_{1-i},R^{+}_{\varepsilon}(X))+\varepsilon$ because $\mbox{because }X_{\varepsilon}^{d^{*}}\subseteq R^{+}_{\varepsilon}(X).$
    So $s_{0}\sim_{\varepsilon}s_{1}$ and hence $\varepsilon\in E^{*}$.
  • $\leq$:
    Let $\varepsilon\in E^{*}$. We prove that for all $\gamma>\varepsilon$, we have $\gamma\in E^{\Delta}$. This will prove that $\inf E^{\Delta}\leq\varepsilon$. Since $\varepsilon$ is arbitrary, this will give the result: $\inf E^{\Delta}\leq\inf E^{*}$.
    So let $\gamma>\varepsilon$. Because $\varepsilon\in E^{*}$, we have $s_{0}\sim_{\varepsilon}s_{1}$, and hence there is some $\varepsilon$-bisimulation $R$ such that $s_{0}Rs_{1}$. We want $\gamma\in E^{\Delta}$, so let $a\in\mathcal{A}$ and $X\subseteq S$. We want $\tau_{a}(s_{i},X)\leq\tau_{a}(s_{1-i},X^{d^{*}}_{\gamma})+\gamma,\quad i=0,1$. First observe that $R(X)\subseteq X^{d^{*}}_{\gamma}$. Indeed, if $y\in R(X)$, then there is some $x\in X$ such that $yRx$. Because $R$ is an $\varepsilon$-bisimulation, $d^{*}(y,x)\leq\varepsilon$, so $y\in X^{d^{*}}_{\gamma}$ (we cannot say $y\in X^{d^{*}}_{\varepsilon}$ because these sets are open balls). Then, for $i=0,1$:
    $\tau_{a}(s_{i},X)$ $\leq\tau_{a}(s_{1-i},R(X))+\varepsilon,$ because $R$ is an $\varepsilon$-bisimulation
    $\leq\tau_{a}(s_{1-i},X^{d^{*}}_{\gamma})+\gamma,$ because $\mbox{because }R(X)\subseteq X^{d^{*}}_{\gamma}\mbox{and }\varepsilon<\gamma.$
    So, $\gamma\in E^{\Delta}$, as wanted.
So we have proven that $d^{*}$ is a fixpoint of $\Delta_{LP}$, and it is the greatest by Proposition 3.9.
At this point we remark again that even if it seems not surprising that the Lévy-Prokhorov distance on distributions has some link with the $\varepsilon$-distance, going from the basic distance on distributions $\delta_{LP}$ to a meaningful distance on LMP is not direct (one has to find the fixed-point properties). Moreover, as far as we know, this is the first time that a distance on distributions other than the Kantorovich distance is used for characterizing behavioural distances.
3.3 The Lévy-Prokhorov lifting of the subdistribution functor to $\mathbf{PMet_{1}}$
Using the LP distance lifting, we define a functor $\mathcal{D}_{LP}$ on $\mathbf{PMet_{1}}$, the category of 1-bounded pseudometric spaces with nonexpansive functions, as follows: On objects $\langle X,d\rangle$, we have
$\mathcal{D}_{LP}\langle X,d\rangle=(\mathcal{D}X,\delta^{d}_{LP})$
and on morphisms $f\colon\langle X_{0},d_{0}\rangle\to\langle X_{1},d_{1}\rangle$ we set $\mathcal{D}_{LP}f=\mathcal{D}f$, that is
$\mathcal{D}_{LP}f\colon\mathcal{D}_{LP}\langle X_{0},d_{0}\rangle\to\mathcal{D}_{LP}\langle X_{1},d_{1}\rangle\textrm{ with }\varphi\mapsto\lambda y.\,\varphi(f^{-1}(\{y\})).$
Proposition 3.18.
$\mathcal{D}_{LP}$ is a functor on $\mathbf{PMet_{1}}$.
Proof 3.19.
The only thing to prove is that $\mathcal{D}_{LP}f$ is nonexpansive. Let $\varphi_{0},\varphi_{1}\in\mathcal{D}_{LP}\langle X_{0},d_{0}\rangle$. Then (we omit the symmetric inequality for simplicity)
$\delta^{d_{1}}_{LP}(\mathcal{D}_{LP}f(\varphi_{0}),\mathcal{D}_{LP}f(\varphi_{1}))$ $=$ $\inf\{\varepsilon\mid\forall B\subseteq X_{1}:\varphi_{0}(f^{-1}(B))\leq\varphi_{1}(f^{-1}(B_{\varepsilon}^{d_{1}}))+\varepsilon\}$
$\stackrel{{\scriptstyle(*)}}{{\leq}}$ $\inf\{\varepsilon\mid\forall B\subseteq X_{1}:\varphi_{0}(f^{-1}(B))\leq\varphi_{1}((f^{-1}B)_{\varepsilon}^{d_{0}})+\varepsilon\}$
$=$ $\inf\{\varepsilon\mid\forall A\in\{f^{-1}(B)\mid B\subseteq X_{1}\}:\varphi_{0}(A)\leq\varphi_{1}(A_{\varepsilon}^{d_{0}})+\varepsilon\}$
$\stackrel{{\scriptstyle(**)}}{{\leq}}$ $\inf\{\varepsilon\mid\forall A\subseteq X_{0}:\varphi_{0}(A)\leq\varphi_{1}(A_{\varepsilon}^{d_{0}})+\varepsilon\}$
$=$ $\delta^{d_{0}}_{LP}(\varphi_{0},\varphi_{1})$
where the marked inequalities are justified below. For the first one, marked with ()(*)( ∗ ), let
 $\mbox{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }V=\{\varepsilon\mid\forall B\subseteq X_{1}:\varphi_{0}(f^{-1}(B))\leq\varphi_{1}((f^{-1}B)_{\varepsilon}^{d_{0}})+\varepsilon\}\mbox{ and}$
$W=\{\varepsilon\mid\forall B\subseteq X_{1}:\varphi_{0}(f^{-1}(B))\leq\varphi_{1}(f^{-1}(B_{\varepsilon}^{d_{1}}))+\varepsilon\}.$
We show that $V\subseteq W$. Let $\varepsilon\in V$. For every $B\subseteq X_{1}$, let us first show that $f^{-1}(B)_{\varepsilon}^{d_{0}}\subseteq f^{-1}(B_{\varepsilon}^{d_{1}})$. Indeed, take $x\in f^{-1}(B)_{\varepsilon}^{d_{0}}$ and let $a\in f^{-1}(B)$ be such that $d_{0}(x,a)<\varepsilon$. Then by nonexpansivity of $f$,
$d_{1}(f(x),f(a))\leq d_{0}(x,a)<\varepsilon$
and so $f(x)\in B_{\varepsilon}^{d_{1}}$, because $f(a)\in B$. So $x\in f^{-1}(B_{\varepsilon}^{d_{1}})$. Now combining this with the inequality condition in $V$ we get
$\varphi_{0}(f^{-1}(B))\leq\varphi_{1}(f^{-1}(B)_{\varepsilon}^{d_{0}})+\varepsilon\leq\varphi_{1}(f^{-1}(B_{\varepsilon}^{d_{1}}))+\varepsilon$
and so $\varepsilon\in W$.
For the inequality marked with (**) , observe that if $\varepsilon$ satisfies the inequality for all $A\subseteq X_{0}$ then it does for all $A\in\{f^{-1}(B)\mid B\subseteq X_{1}\}$. So the infimum is taken on a possibly bigger set on the left-hand side of the inequality, which concludes the argument.
Just like for the Kantorovich-lifting of the distribution functor, we can prove that the Lévy-Prokhorov lifting of the distribution functor is locally nonexpansive, in the next proposition. For this reason, note that given pseudometric spaces $\langle X,d_{X}\rangle$ and $\langle Y,d_{Y}\rangle$ in $\mathbf{PMet_{1}}$, the hom-set $X\to Y$ of all nonexpansive maps from $X$ to $Y$ carries a metric defined by
$d_{X\to Y}(f_{1},f_{2})=\sup_{x\in X}d_{Y}(f_{1}(x),f_{2}(x)).$
Proposition 3.20.
The functor $\mathcal{D}_{LP}$ is locally nonexpansive, that is, for $f_{1},f_{2}\in X\to Y$
$\delta_{\mathcal{D}_{LP}X\to\mathcal{D}_{LP}Y}(\mathcal{D}_{LP}f_{1},\mathcal{D}_{LP}f_{2})\leq d_{X\to Y}(f_{1},f_{2}).$
Proof 3.21.
Let $f_{1},f_{2}\in X\to Y$. We have that
$\delta_{\mathcal{D}_{LP}X\to\mathcal{D}_{LP}Y}(\mathcal{D}_{LP}f_{1},\mathcal{D}_{LP}f_{2})=\sup_{\varphi\in\mathcal{D}X}\delta_{LP}^{d_{Y}}(\mathcal{D}_{LP}f_{1}(\varphi),\mathcal{D}_{LP}f_{2}(\varphi)).$
Let $\varphi\in\mathcal{D}X$ and recall that $\mathcal{D}_{LP}f_{i}(\varphi)=\varphi(f_{i}^{-1}(\cdot))$. We will show that
$\delta_{LP}^{d_{Y}}(\varphi(f_{1}^{-1}(\cdot)),\varphi(f_{2}^{-1}(\cdot)))\leq d_{X\to Y}(f_{1},f_{2}).$
Let $\alpha=d_{X\to Y}(f_{1},f_{2})=\sup_{x\in X}d_{Y}(f_{1}(x),f_{2}(x))$. Let $B\subseteq Y$. We show $f_{1}^{-1}(B)\subseteq f_{2}^{-1}(B^{d_{Y}}_{\alpha})$
$f_{1}^{-1}(B)$ =$=\{x\mid\exists y\in B:y=f_{1}(x)\}$
=$=\{x\mid\exists y\in B:y=f_{1}(x)\mbox{ and }d_{Y}(y,f_{2}(x))\leq\alpha\}$ since $\mbox{since }d_{Y}(f_{1}(x),f_{2}(x))\leq\alpha$
$\subseteq\{x\mid\exists y\in B:d_{Y}(y,f_{2}(x))\leq\alpha\}$
=$=\{x\mid f_{2}(x)\in B^{d_{Y}}_{\alpha}\}$
$=f_{2}^{-1}(B^{d_{Y}}_{\alpha}).$
This implies that $\alpha$ satisfies $(\forall B\subseteq Y:\varphi(f_{1}^{-1}(B))\leq\varphi(f_{2}^{-1}(B^{d_{Y}}_{\alpha}))+\alpha)$, and similarly with 1 and 2 inverted.
$\delta_{LP}^{d_{Y}}(\varphi(f_{1}^{-1}(\cdot)),\varphi(f_{2}^{-1}(\cdot)))$ $=\inf\{\varepsilon\mid\forall B\subseteq Y:\varphi(f_{1}^{-1}(B))\leq\varphi(f_{2}^{-1}(B^{d_{Y}}_{\varepsilon}))+\varepsilon\}$
$\leq\alpha\,\,=\,\,d_{X\to Y}(f_{1},f_{2}),$
as wanted.
3.4 The lifted functor $\mathcal{D}_{LP}$ is not a monad lifting of $\mathcal{D}$
Behavioural distances have been axiomatized within the line of work on quantitative equational theories [38]. It was therefore a natural question for us whether the $\varepsilon$-distance, or the Lévy-Prokhorov distance itself can be given a quantitative axiomatization. This is what we briefly investigate in this section.
Mio et al. [42, Lemma 7.2, Theorem 7.7(2)] have shown that:
  • If a monad on metric spaces is axiomatizable, then (just by being a monad on metric spaces) it has nonexpansive unit and multiplication.
  • A monad on metric spaces that is a lifting of a monad on $\mathbf{Sets}$ is axiomatizable with a quantitative theory. Being a lifting means that it acts on objects and arrows in metric spaces in the same way as it does in $\mathbf{Sets}$ and that the unit and multiplication are nonexpansive.
The Kantorovich lifting of $\mathcal{D}$ has been axiomatized, and the total variation distance has been shown non-axiomatizable (as the unit is not nonexpansive). In the case of $\mathcal{D}_{LP}$ we can see that the unit is nonexpansive, but the multiplication is not, and hence $\mathcal{D}_{LP}$ is not a lifting of the monad $\mathcal{D}$ to $\mathbf{PMet_{1}}$ and the Lévy-Prokhorov distance on distributions is not axiomatizable with a quantitative theory, at least not with the standard multiplication.
Recall the definitions of $\eta$ and $\mu$ for the subdistribution monad $\mathcal{D}$:
$\eta_{X}(x)=1_{x},\text{ the Dirac distribution at }x;\text{ and }\mu_{X}(\Phi)(x)=\sum_{\varphi}\Phi(\varphi)\cdot\varphi(x).$
Lemma 3.22.
The unit $\eta$ of $\mathcal{D}$ is nonexpansive with respect to the Lévy-Prokhorov distance. Moreover, it is an isometry, i.e., $\delta_{LP}^{d}(1_{x},1_{y})=d(x,y)$.
Proof 3.23.
We have
$\delta^{d}_{LP}(\eta_{X}(x),\eta_{X}(y))$ $=\delta^{d}_{LP}(1_{x},1_{y})$
$=\inf\{\varepsilon\mid\forall A\subseteq X:1_{x}(A)\leq 1_{y}(A^{d}_{\varepsilon})+\varepsilon,1_{y}(A)\leq 1_{x}(A^{d}_{\varepsilon})+\varepsilon\}.$
Let $A$ be a subset of $X$. We distinguish four cases in which we consider the relevant inequalities:
$1_{x}(A)$ $\leq 1_{y}(A^{d}_{\varepsilon})+\varepsilon$ (7)
[…]
  1. 1.
    $x\notin A,y\notin A$: Then $01_{x}(A)=$ 0 and $01_{y}(A)=$ 0, and any $\varepsilon\geq 0$ satisfies both (7) and (8).
  2. 2.
    $x\notin A,y\in A$: Here $01_{x}(A)=$ 0 and $11_{y}(A)=$ 1, and  (7) holds always, but (8) holds for $\varepsilon>d(x,y)$ as then $x\in A^{d}_{\varepsilon}$ and hence $11_{x}(A^{d}_{\varepsilon})=$ 1, and might not hold for $\varepsilon\leq d(x,y)$, e.g. when $A=\{y\}$ and $\varepsilon<d(x,y)\leq 1$, as then $01_{x}(A^{d}_{\varepsilon})=$ 0.
  3. 3.
    $x\in A,y\notin A$: This case is dual to the second case. Here (8) always holds, but (7) holds for $\varepsilon>d(x,y)$ and might not hold for $\varepsilon\leq d(x,y)$ as, e.g., for $A=\{x\}$ we have $y\in A^{d}_{\varepsilon}$ if and only if $\varepsilon>d(x,y)$.
  4. 4.
    $x\in A,y\in A$: Then $11_{x}(A)=$ 1 and $11_{y}(A)=$ 1, as well as $11_{x}(A^{d}_{\varepsilon})=1_{y}(A^{d}_{\varepsilon})=$ 1, and again, as in the first case, any $\varepsilon\geq 0$ satisfies both (7) and (8).
Hence, both inequalities are satisfied for all $A$ iff $\varepsilon>d(x,y)$, and therefore $\delta^{d}_{LP}(1_{x},1_{y})=d(x,y)$.
The unit of the distribution monad is also an isometry with respect to the Kantorovich distance (without discount), as is easy to prove using the duality with the Wasserstein distance.
However, the multiplication $\mu$ of the monad $\mathcal{D}$ is not nonexpansive, as the following example shows.
Example 3.24.
[…]
The ambient distance is $d(\bot,\bullet)=1$, we omit it from the notation: we write $\delta_{LP}$ for $\delta_{LP}^{d}$. Consider the distributions $\varphi:=1_{a}$ and $\psi:=(1-\varepsilon)1_{b}+\varepsilon 1_{c}$ over these distributions, as pictured. We collect several useful facts:
  • $\delta_{LP}(a,b)=\varepsilon$ and $\delta_{LP}^{\delta_{LP}}(\varphi,\psi)=\varepsilon$.
  • $\mu\varphi=a$ and $\mu\psi=(1-\varepsilon)^{2}\bot+(1-\varepsilon)\varepsilon\bullet+\varepsilon\bullet=(1-\varepsilon)^{2}\bot+(2-\varepsilon)\varepsilon\bullet$.
  • {}$\{\bot\}_{\gamma}=\{\bot\}$ for any $\gamma\leq\varepsilon$ and hence $\mu\varphi(\{\bot\})=1$ and $\mu\psi(\{\bot\}_{\gamma})=\mu\psi(\{\bot\})=(1-\varepsilon)^{2}$.
For $\gamma<\varepsilon(2-\varepsilon)$, we have $1\not\leq(1-\varepsilon)^{2}+\gamma$ which therefore yields
$\mu\varphi(\{\bot\})=1\not\leq(1-\varepsilon)^{2}+\gamma=\mu\psi(\{\bot\}_{\gamma})+\gamma.$
This shows that $\delta_{LP}(\mu\varphi,\mu\psi)\geq\varepsilon(2-\varepsilon)>\varepsilon=\delta_{LP}^{\delta_{LP}}(\varphi,\psi)$ and hence $\mu$ is not nonexpansive.
This result is not surprising, as the binding operator $\mu$ really is a multiplication, and so it accords well with the Kantorovich metric, which does multiply the probabilities of distributions in play. We do not know yet if another operator could help express the functor $\mathcal{D}_{LP}$ as a monad lifting.
4 $\varepsilon$-(Bi)simulations, coalgebraically
The Kantorovich behavioural distance has a coalgebraic characterization via a final coalgebra semantics. Whether the $\varepsilon$-distance is obtained by finality is still open, but the results on the Kantorovich distance do not seem to directly apply here. However, as the distance is defined via $\varepsilon$-bisimulations and bisimilarity, it is natural to see whether a coalgebraic semantics arising from generalizing coalgebraic bisimulations and bisimilarity is in place. The answer to this question is positive and we present the necessary observations in this section. An abstract (in this case coalgebraic) characterization allows for, on the theoretical side, deeper and clearer understanding of the notion under study, and, on the practical side, generalizations. We could use this generality to define notions of $\varepsilon$-bisimulations for other types of (probabilistic) systems.
In a nutshell, in this section we show that $\varepsilon$-simulation and $\varepsilon$-bisimulation have a span-diagram characterization, in a way similar to coalgebraic simulation [33, 31, 32] and Aczel-Mendler coalgebraic bisimulation [3, 34, 45]. The development is using a notion of $\varepsilon$-coupling defined in [51]. For simplicity, we ignore the labels in this section, i.e., we assume there is a single label.
To start with, we recall the basic notions related to coalgebraic (bi)simulation, formulated for the functor $\mathcal{D}$ that we are mainly interested in in this work. Recall the definition of the sub-distribution functor $\mathcal{D}$ on $\mathbf{Sets}$: It maps a set $X$ to $\mathcal{D}X$, the set of discrete subdistributions on $X$ and on arrows $f\colon X\to Y$, $\mathcal{D}f\colon\mathcal{D}X\to\mathcal{D}Y$ is defined by
$\mathcal{D}f(\varphi)(y)=\sum_{x\in f^{-1}(\{y\})}\varphi(x)=\varphi(f^{-1}(\{y\}).$
Coalgebras of the functor $\mathcal{D}$ are Markov chains, formally they are pairs $(X,c)$ of a carrier set $X$ and transition map $c\colon X\to\mathcal{D}X$. We often just refer to the coalgebra by the transition map $c$. Given two such coalgebras, $c\colon X\to\mathcal{D}X$ and $d\colon Y\to\mathcal{D}Y$ a coalgebra homomorphism from $(X,c)$ to $(Y,d)$ is a map $h\colon X\to Y$ making the following diagram commute:
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{h}$$\scriptstyle{c}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d}$${\mathcal{D}X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{D}h}$${\mathcal{D}Y}$
Definition 4.1.
Given two coalgebras $c\colon X\to\mathcal{D}X$ and $d\colon Y\to\mathcal{D}Y$, a coalgebraic bisimulation is a relation $R\subseteq X\times Y$ such that there exists a coalgebra structure $b\colon R\to\mathcal{D}R$ making the two projections $\pi_{1}\colon R\to X$ and $\pi_{2}\colon R\to Y$ coalgebra homomorphisms, i.e., making the following span diagram commute:
${X\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{c}$${R\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi_{1}}$$\scriptstyle{\pi_{2}}$$\scriptstyle{b}$${Y\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{d}$${\mathcal{D}X}$${\mathcal{D}R\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathcal{D}\pi_{1}}$$\scriptstyle{\mathcal{D}\pi_{2}}$${\mathcal{D}Y}$
It is well known, see [20, 47] that this notion of bisimulation corresponds to the standard notion of probabilistic bisimulation due to Larsen and Skou [37]. For the case of continuous distributions, taking a cospan instead of a span helps avoiding the need for analytic spaces [16]. Back to the discrete case, one way to define the coalgebra structure $b$ is using the notion of a coupling: If for all $(x,y)\in R$, there is a coupling $\beta$ for $\mu=c(x)$ and $\nu=d(y)$, then setting $b(x,y)=\beta$ provides the needed transition structure. Recall that $\beta\in\mathcal{D}R$ is a coupling of $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$ if its marginals are $\mu$ and $\nu$, respectively, that is
$\sum_{y\in Y}\beta(x,y)=\mu(x),\quad\sum_{x\in X}\beta(x,y)=\nu(y).$
Note that this definition of a coupling, fitting the definition of coalgebraic bisimulation for the subprobability distribution functor, ensures that a coupling of two subdistributions exists only if they have the same total mass. One could also define couplings by adding a dummy element and hence viewing a subdistribution as a full distribution, which provides a more general notion. However, these details are not relevant for what we are really interested in, which are $\varepsilon$-bisimulations.
We are going to use a relaxed notion of $\varepsilon$-coupling [51] to define $\varepsilon$-(bi)simulation in analogy with the coalgebraic definition above. Before we recall those, let us first mention coalgebraic simulations, due to [33, 31, 32] with a coalgebraic formulation for LMPs already in [21].
For this, note that an ordered functor $F$ is a functor with order on each object $FX$. The subdistribution functor $\mathcal{D}$ is ordered, e.g., by pointwise order. Note that this order becomes trivial in case of the distribution functor which is the reason why we work with subdistributions here. Coalgebraic simulation can be defined for ordered functors, here we only recall the definition in the special case of the subdistribution functor $\mathcal{D}$. All notions involve lax and oplax morphism.
Definition 4.2.
A lax homomorphism from $(X,c)$ to $(Y,d)$ is a morphism $l\colon X\to Y$ with the property that $d\circ l\sqsubseteq\mathcal{D}l\circ c$, i.e., it makes the left lax diagram below commute. An oplax homomorphism from $(X,c)$ to $(Y,d)$ is a lax homomorphism for the dual/opposite order, that is a morphism $o\colon X\to Y$ that makes the right diagram below commute:
[…]
The order on distributions is defined pointwise: For a set $A$ and $\varphi,\psi\in\mathcal{D}A$, we have $\varphi\sqsubseteq\psi$ iff for all $a\in A$, $\varphi(a)\leq\psi(a)$.
Note that, equivalently, $\varphi\sqsubseteq\psi$ iff for all $B\subseteq A$, $\varphi(B)\leq\psi(B)$, where the right-to-left implication follows from instantiating on singleton subsets and the left-to-right implication follows from the additivity of distributions.
Initially, generic coalgebraic simulations have been studied for ordered functors, by Hughes and Jacobs [33], using a notion of lax relation lifting. On the other hand, Hasuo [31, 32] discovered (generalizations of) forward, backward, and hybrid simulations in Kleisli categories and moreover showed in [32] that Hughes-Jacobs simulations are a special case. While Kleisli categories are appealing for soundness results, and monads suitable for traces come equipped with an order, we may safely ignore all Kleisli aspects here. We now recall the following notion of simulations [31], formulated for general categories of coalgebras (as long as the functor comes equipped with an order).
Definition 4.3.
A forward simulation from $(X,c)$ to $(Y,d)$ is a lax homomorphism from $(Y,d)$ to $(X,c)$. A backward simulation from $(X,c)$ to $(Y,d)$ is an oplax homomorphism. Combining the lax-commuting boxes leads to hybrid, forward-backward and backward-forward, simulations. In particular, Hughes-Jacobs simulation is a forward-backward simulation: It is a relation $R$ on which there exists a coalgebra structure $b\colon R\to\mathcal{D}R$ making $\pi_{1}$ a forward simulation from $(X,c)$ to $(R,b)$ and $\pi_{2}$ a backward simulation from $(R,b)$ to $(Y,d)$, depicted:
[…]
Clearly, a symmetric simulation on a single system is a bisimulation.
In the rest of this section, we will show that $\varepsilon$-(bi)simulations can be depicted similarly. For this we will need bounded-lax-commutativity of morphisms as well as a notion of $\varepsilon$-coupling (which then directly gives an $\varepsilon$-relation-lifting). We will focus here on the functor $\mathcal{D}$ only. These notions can be generalized to generic coalgebras for functors with suitable structure. However, those observations are beyond the scope of this paper and we will elaborate on them elsewhere.
Definition 4.4.
An $\varepsilon$-lax homomorphism from $(X,c)$ to $(Y,d)$ is a morphism $l\colon X\to Y$ that makes the left $\varepsilon$-lax diagram below commute, i.e., $d\circ l\sqsubseteq_{\varepsilon}\mathcal{D}l\circ c$. An $\varepsilon$-oplax homomorphism from $(X,c)$ to $(Y,d)$ is a morphism $o\colon X\to Y$ with the property that $\mathcal{D}o\circ c\sqsubseteq_{\varepsilon}d\circ o$, i.e., it makes the right $\varepsilon$-lax diagram below commute.
[…]
The $\varepsilon$-order on distributions is defined by: For a set $A$ and $\varphi,\psi\in\mathcal{D}A$, we have $\varphi\sqsubseteq_{\varepsilon}\psi$ iff for all $B\subseteq A$, $\varphi(B)\leq\psi(B)+\varepsilon$.
Remark 4.5.
Note that $\sqsubseteq_{\varepsilon}$ is not an order relation, namely it is not transitive. Note also that the definition of $\sqsubseteq_{\varepsilon}$ can not be expressed on elements (of $A$) only, as it is stronger than the property: for all $a\in A$, $\varphi(a)\leq\psi(a)+\varepsilon$. Namely, the role of $\varepsilon$ here, just like in the definition of $\varepsilon$-(bi)simulation is global.
We next recall the notion of $\varepsilon$-coupling from [51].
Definition 4.6.
Let $R\subseteq X\times Y$. A distribution $\beta\in\mathcal{D}R$ is an $\varepsilon$-coupling for $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$ iff the following three conditions hold:
  • 1.
    $\sum_{y\in Y}\beta(x,y)\leq\mu(x)$, for all $x\in X$
  • 2.
    $\sum_{x\in X}\beta(x,y)\leq\nu(y)$, for all $y\in Y$
  • 3.
    $\mu(X)\leq\sum_{x\in X,y\in Y}\beta(x,y)+\varepsilon=\sum_{(x,y)\in R}\beta(x,y)+\varepsilon$.
If we want to be fully precise, we should consider that $\beta\in\mathcal{D}(X\times Y)$ with support contained in $R$. We trust the readers can tolerate the freedom we took to write $\beta\in\mathcal{D}R$ and yet sometimes sum over all $x\in X,y\in Y$, i.e., identify $\beta\in\mathcal{D}R$ with a distribution in $\mathcal{D}(X\times Y)$ that assigns probability 00 to all pairs out of $R$ and acts as $\beta$ on $R$.
Before we proceed with the main observation of this section, we prove an auxiliary property that allows for rewriting the definition of $\varepsilon$-coupling.
Lemma 4.7.
Assume that Condition 1. in Definition 4.6 holds for $\beta\in\mathcal{D}R$ with $R\subseteq X\times Y$ and $\mu\in\mathcal{D}X$. Then $\mu(X)\leq\sum_{x\in X,y\in Y}\beta(x,y)+\varepsilon$ (i.e., Condition 3. in Definition 4.6) is equivalent to the condition
$\mu(S)\leq\sum_{x\in S,y\in Y}\beta(x,y)+\varepsilon,\quad\text{for all }S\subseteq X.$
Proof 4.8.
The right-to-left implication is immediate, as $X\subseteq X$. For the left-to-right implication, assume that for some $S\subseteq X$ we have
$\mu(S)>\sum_{x\in S,y\in Y}\beta(x,y)+\varepsilon.$
Since, by assumption, for all $x\in X$, and hence for all $x\in X\setminus S$, $\mu(x)\geq\sum_{y\in Y}\beta(x,y)$, we have
$\mu(X)$ $=$ $\mu(S)+\mu(X\setminus S)$
$>$ $\sum_{x\in S,y\in Y}\beta(x,y)+\varepsilon+\sum_{x\in X\setminus S}\mu(x)$
$\geq$ $\sum_{x\in S,y\in Y}\beta(x,y)+\sum_{x\in X\setminus S}\sum_{y\in Y}\beta(x,y)+\varepsilon$
$=$ $\sum_{x\in X,y\in Y}\beta(x,y)+\varepsilon.$
The property now follows by contraposition.
This property again emphasises the global nature of $\varepsilon$ in our situation. As a consequence, for an $\varepsilon$-coupling $\beta\in\mathcal{D}R$ of $\mu\in\mathcal{D}X$ and $\nu\in\mathcal{D}Y$, with $R\subseteq X\times Y$, for each $S\subseteq X$:
$\sum_{x\in S,y\in Y}\beta(x,y)\,\,\leq\,\,\mu(S)\,\,\leq\,\,\sum_{x\in S,y\in Y}\beta(x,y)+\varepsilon.$
Proposition 4.9.
Let $(X,c)$ and $(Y,d)$ be two $\mathcal{D}$-coalgebras. The following three properties are equivalent for $R\subseteq X\times Y$:
  • 1.
    $R$ is an $\varepsilon$-simulation.
  • 2.
    For every $(x,y)\in R$, there is an $\varepsilon$-coupling $\beta\in\mathcal{D}R$ of $\mu=c(x)$ and $\nu=d(y)$.
  • 3.
    The following "$\varepsilon$-lax-bounded" span of morphisms commutes:
[…]
Proof 4.10.
The equivalence of 1. and 2. has been shown in [51]. Let $R$ be an $\varepsilon$-simulation. For $(x,y)\in R$, let $\mu=c(x)$ and $\nu=d(y)$. Define $b\colon R\to\mathcal{D}R$ by $(x,y)\mapsto\beta$, the epsilon coupling of $\mu$ and $\nu$ that exists.
Then, unfolding the definitions, we get:
  • $\mathcal{D}\pi_{1}\circ b\sqsubseteq c\circ\pi_{1}$ is equivalent to: for all $(x,y)\in R$, condition 1. from Definition 4.6 holds.
  • $\mathcal{D}\pi_{2}\circ b\sqsubseteq d\circ\pi_{2}$ is equivalent to: for all $(x,y)\in R$, condition 2. from Definition 4.6 holds.
  • $c\circ\pi_{1}\sqsubseteq_{\varepsilon}\mathcal{D}\pi_{1}\circ b$ is equivalent to: for all $(x,y)\in R$, for all $S\subseteq X$,
    $\mu(S)\leq\sum_{x\in S,y\in Y}\beta(x,y)+\varepsilon$
    which by Lemma 4.7 is equivalent to condition 3. from Definition 4.6.
These facts yield the equivalence of 3. with 2. (and hence 1. as well).
Note that Condition 2. in Proposition 4.9 can be taken as a definition of $\varepsilon$-relation-lifting, in analogy to relation lifting for $\mathcal{D}$ being defined using the existence of a coupling for any pair of elements in $R$. This way was also taken in [33] with the definition of lax relation lifting.
As a consequence of Proposition 4.9, we immediately get that a relation $R\subseteq X\times X$ on the states of a $\mathcal{D}$-coalgebras is an $\varepsilon$-bisimulation iff the following diagram commutes:
[…]
As already mentioned, generalizing the notions of approximate (bi)simulations to generic coalgebras is an interesting direction that we will undertake in the near future. A similar notion is being developed for cost automata [19].
5 Concluding remarks
We have shown that $\varepsilon$-bisimulations are closely connected with the Lévy-Prokhorov pseudometric on (sub)probability distributions: The LP pseudometric lifts $\mathcal{D}$ to a functor on pseudometric spaces and induces coinductive behaviour distance on LMPs, which is the greatest fixpoint of a suitable functional and turns out to be exactly the distance induced by $\varepsilon$-bisimilarity. Remarkably, any fixpoint distance of that functional defines an $\varepsilon$-bisimulation. This is the first time that a distance on distributions other than the Kantorovich distance is used for characterizing behavioural distances.
6 Acknowledgement
This work was partly done during a sabbatical of Josée Desharnais at the University of Salzburg, as well as during Dagstuhl Seminar 24432 and the Bellairs Workshop on Quantitative Reasoning 2025. We thank these venues, the organizers, and the participants for providing a perfect working environment. We also thank Matteo Mio and Franck van Breugel for some fruitful and motivating discussions.